# A Smooth Transition from Giant Planets to Brown Dwarfs from the Radial Occurrence Distribution

### v1

@@ 3.1 Joint RV/astrometry orbital fits
Figure 1: Left: The masses and semi-major axes of the substellar ($M_{c}\sin i\leq 80\,M_{\mathrm{Jup}}$) companions reported in the CLS catalog. Orange points show the average $a$ and $M_{c}\sin i$ fitted, or in some cases quoted, by CLS, while blue points show the average $a$ and $M_{c}$ refitted in this work. Green borders indicate companions in whose orbital fits we included HGCA astrometry. The color map shows the CLS sensitivity to true mass, averaged over all 719 stars in the survey, with contours indicating sensitivity deciles and the 50% completeness line marked in black. Horizontal green lines mark 13 and 80 $M_{\mathrm{Jup}}$, the traditional brown dwarf mass limits. Right: Our re-fit catalog of 195 CLS companions. Contours enclose 68% of an individual companion’s posterior draws. Black rectangles indicate the cells in which we calculated companion occurrence (see Section 4). The completeness map is the same as in the right panel. Note that because the colored markers in the left panel indicate averages, companions with $M_{c}>80\,M_{\mathrm{Jup}}$ may still have significant posterior mass below this limit, and vice versa.
3.2 Occurrence Framework
@@
We modeled planet occurrence as a piecewise constant function, that is, a histogram, in companion mass $M_{c}$ and semi-major axis $a$. We assumed that the companion occurrence rate in each histogram bin was independent of those in other bins.
@@
Following previous studies (e.g., ForemanMackey2014; Fulton2021; VanZandt2025), we modeled our catalog as a series of draws from a censored Poisson point process defined by the companion occurrence rate density (ORD) $\lambda(\log a,\log M_{c}|\boldsymbol{\theta})$, which gives the number of companions per star per logarithmic $a$-$M_{c}$ interval. We use $\boldsymbol{\theta}$ to denote the vector of model parameters of $\lambda$, which for a histogram model is simply the bin heights.
Our full likelihood is given by
@@
$\mathcal{L}\propto e^{-\Lambda}\prod_{n=1}^{N_{c}}\lambda(\boldsymbol{\omega}_{n}|\boldsymbol{\theta})Q(\boldsymbol{\omega}_{n}),$ (1)
where we define $\boldsymbol{\omega}=(\log a,\log M_{c})$ for notational compactness. This likelihood function comprises two terms. The first gives the probability of $N_{c}$ companion detections over the two-dimensional parameter space:
$P(N_{c}|\boldsymbol{\theta})=\frac{e^{-\Lambda}\Lambda^{N_{c}}}{N_{c}!}.$ (2)
@@
The Poisson rate parameter $\Lambda$ gives the expected total number of observed companions among $N_{*}$ host stars in the interval of interest:
$\Lambda=N_{*}\int\lambda(\boldsymbol{\omega}|\boldsymbol{\theta})\,Q(\boldsymbol{\omega})\,d\boldsymbol{\omega},$ (3)
@@
where $Q$, the average sensitivity of the CLS to companions as a function of separation and true mass, accounts for missed companions.
@@
The second term gives the probability that the $N_{c}$ detected companions possess the set of observed parameters $\{\boldsymbol{\omega}\}$:
@@
$P(\{\boldsymbol{\omega}\}|N_{c},\boldsymbol{\theta})=\prod^{N_{c}}_{n=1}\frac{\lambda(\boldsymbol{\omega}_{n}|\boldsymbol{\theta})Q(\boldsymbol{\omega}_{n})}{\Lambda}.$ (4)
We employed Equation 1 for each bin individually, treating the interval defined by the bin as the parameter space of interest and $\boldsymbol{\theta}$ as a single number giving the ORD in that bin.
Equation 1 gives the model likelihood in the case of negligible $a$ and $M_{c}$ uncertainties. We accounted for our non-negligible uncertainties as follows. We drew one sample from each companion posterior in our catalog to obtain a “catalog draw,” for which we then inferred a distribution of ORD values by evaluating Equation 1 over a fine grid in $\lambda$. We repeated this procedure 1000 times and averaged the resulting distributions to marginalize over our uncertainty in the true companion parameters. Our catalog sampling method ensured that detected companions influenced the inferred occurrence rate in a given bin in proportion to the fraction of their posterior mass that overlapped that bin.
4 Results
4.1 Companion occurrence as a function of mass and semi-major axis
@@
We measured occurrence as a function of both mass and separation in 30 cells defined by the following boundaries: $\frac{M_{c}}{M_{\mathrm{Jup}}}=${0.8, 1.6, 3.2, 6.4, 13, 80}, $\frac{a}{AU}=${0.03, 0.1, 0.3, 1, 3, 10, 30}. We highlight our region of interest in the context of the CLS true mass catalog in the right panel of Figure 1. Figure 2 shows the results of these calculations. The drawback of calculating occurrence in 30 separate bin is that each bin contains few companion detections, yielding large uncertainties on our measurements. We therefore performed a simple model comparison for the semi-major axis distribution of each mass bin: a step function versus a log-linear model, given respectively by
@@
4.2 The Brown Dwarf Desert at 1-10 AU
@@
Figure 3: Marginalized occurrence rate density of CLS companions between 1–10 AU. Companion occurrence decreases log-linearly with increasing mass up to $\sim$20 $M_{\mathrm{Jup}}$, disfavoring a ‘cut-off’ mass at which companion formation near the ice line abruptly becomes inefficient. We annotate each mass bin with the true occurrence rate, quoting a single value for the range 13–80 $M_{\mathrm{Jup}}$ to facilitate comparison with other studies.
4.3 A Rise in Giant Companion Occurrence at the Ice Line
Finally, we examined the occurrence distribution of giant companions as a function of semi-major axis. We used the same separation bins as in Section 4.1, this time marginalizing over companion masses between 0.8–80 $M_{\mathrm{Jup}}$. We found that companion occurrence is approximately constant at close separations, increases abruptly near 1 AU, and remains enhanced to around 10 AU. There is also a marginally significant decrease in occurrence beyond 10 AU, though the uncertainty on our occurrence measurement between 10–30 AU prevents a firm conclusion. Figure 4 shows our marginalized semi-major axis distribution. Comparing this distribution with those we calculated for separate mass intervals (see Section 4.1), we see that the relatively high occurrence between 1–10 AU exhibited in the low-mass bins contributes to the rise at 1 AU in the marginalized distribution. Meanwhile, the small number of detections between 10–30 AU in any mass interval results in comparatively low occurrence in this range. We fit this distribution with both a two-level step function and a broken power law. We found that both models performed comparably well, and that the step location reliably fell near 1 AU.
We repeated these calculations using companion minimum mass and modified parameter bounds to match the analysis of Fulton2021, who calculated the semi-major axis distribution between 0.1–30 AU for companions with $M_{c}\sin i$ =30–6000 $M_{\oplus}$ (0.09–18.9 $M_{\mathrm{Jup}}$). The maximum likelihood parameters of our broken power law fit to this occurrence distribution agreed closely with those derived by Fulton2021, and we recovered the two key features noted in that study: an abrupt rise in occurrence near 1 AU and a possible fall-off beyond 10 AU. We obtained similar results after repeating our calculation for true mass instead of $M_{c}\sin i$. The persistence of the increase in occurrence near 1 AU — albeit at marginal significance — through changes in binning, mass limits, and $M_{c}$ versus $M_{c}\sin i$ suggests that it is a robust feature of the radial occurrence distribution.
@@ 4.3 A Rise in Giant Companion Occurrence at the Ice Line
5 Discussion
5.1 An Astrophysically Motivated Planet-Brown Dwarf Divide
@@
The International Astronomical Union defined brown dwarfs in 2003 as “substellar objects with true masses above the limiting mass for thermonuclear fusion of deuterium $\sim$13 $M_{\mathrm{Jup}}$ ] … no matter how they formed nor where they are located.” [^2] https://web.archive.org/web/20141216075559/http://home.dtm.ciw.edu/users/boss/definition.html This categorization had already been questioned by that time (Burrows2001), and has faced further scrutiny in the years since (e.g., Spiegel2011). Despite the benefit of providing an objective dividing line based on the measurable quantity $M_{c}$, there is no evidence that the processes that form giant planets and brown dwarfs are directly tied to the deuterium burning limit. To better address the nuances of companion formation at the planet-brown dwarf transition, astronomers have searched for definitions motivated by factors such as formation pathway, dynamics, and host star properties. Schlaufman2018 found evidence for a divide between objects above/below 4–10 $M_{\mathrm{Jup}}$ based on the preference for objects below this range to orbit metal-rich host stars, a trait not exhibited by objects above it. Combining high-contrast imaging and astrometric observations between 5–100 AU, Bowler2020 found that companions with $M_{c}<15$ $M_{\mathrm{Jup}}$ have preferentially low eccentricities, peaking at $e\lesssim 0.2$, whereas higher-mass companions exhibit a broad range of eccentricities between ee=0.6–0.9. Using a subset of the companion sample of Bowler2020 but a different framework for calculating parameter priors, DoO2023 found insufficient evidence for a distinction in the eccentricity distributions between low- and high-mass companions. Recent work by our team (Giacalone et al., in press; Gilbert et al., in press) examines the role of both metallicity and eccentricity in the planet-BD distinction in future work.
@@
We searched for a similar distinction in the radial occurrence distributions of planets and BDs in Section 4.1. Among the five mass bins we analyzed, we did not find any abrupt change in the radial distributions with increasing mass. The lack of a sharp dividing line in mass among these distributions suggests that “bottom-up” and “top-down” formation mechanisms may produce companions in overlapping mass ranges. Because the majority of our cells contained fewer than five objects in this experiment, we were unable to constrain occurrence at high significance. Thus, some latent separation-dependent trends may exist that will provide a clearer dividing line between planets and brown dwarfs. On the other hand, we also found smooth variation with increasing mass for the marginalized occurrence distribution in Section 4.2, supporting the idea of a “fuzzy” planet-BD boundary.
@@
We found marginal evidence for jumps in occurrence near 1 AU for companions between 0.8–3.2 $M_{\mathrm{Jup}}$. These features likely drive the analogous and more significant jump in the marginalized radial distribution (Section 4.3). Increased core accretion efficiency is expected near the ice line due to the enhancement of icy pebbles and silicate dust that can coagulate via gravitational and/or streaming instabilities in this region (DrazkowskaAlibert2017; Hyodo2019). Thus, the presence or absence of an abrupt increase in occurrence near the ice line may be used to infer whether an object formed by core accretion or gas/disk instability. Increased Doppler baselines, larger stellar samples, and superior orbital fits through the inclusion of Gaia DR4 astrometry (Winn2022) will enable occurrence measurements precise enough to confirm or rule out an occurrence jump in narrow mass intervals such as those examined here. Such measurements will place an astrophysically motivated planet-brown dwarf dividing line within reach.
5.2 The Brown Dwarf Desert

### v2

@@ III.1 Joint RV/astrometry orbital fits
Figure 1: Top: The masses and semi-major axes of the substellar ($M_{c}\sin i\leq 80\,M_{\mathrm{Jup}}$) companions reported in the CLS catalog. Orange points show the average $a$ and $M_{c}\sin i$ fitted, or in some cases quoted, by CLS, while blue points show the average $a$ and $M_{c}$ refitted in this work. Green borders indicate companions in whose orbital fits we included HGCA astrometry. The color map shows the CLS sensitivity to true mass, averaged over all 719 stars in the survey, with contours indicating sensitivity deciles and the 50% completeness line marked in black. Horizontal green lines mark 13 and 80 $M_{\mathrm{Jup}}$, the traditional brown dwarf mass limits. Bottom: Our re-fit catalog of 194 CLS companions. Contours enclose 68% of an individual companion’s posterior draws. Black rectangles indicate the cells in which we calculated companion occurrence (see Section IV). Our occurrence results flow from the companions whose posteriors fall fully or partially within these rectangles. The completeness map is the same as in the top panel. Note that because the colored markers in the top panel indicate averages, companions with $M_{c}>80\,M_{\mathrm{Jup}}$ may still have significant posterior mass below this limit, and vice versa. We highlight some of the outliers in this plot in Figure 6.
III.2 Occurrence Framework
@@
We modeled companion occurrence as a piecewise constant function, that is, a histogram, in companion mass $M_{c}$ and semi-major axis $a$. We assumed that the companion occurrence rate in each histogram bin was independent of those in other bins.
@@
Following previous studies (e.g., Foreman-Mackey et al. 2014; Fulton et al. 2021; Rogers and Owen 2021; Van Zandt et al. 2025), we modeled our companion catalog as a series of draws from a censored Poisson point process defined by the companion occurrence rate density (ORD) $\lambda(\log a,\log M_{c}|\boldsymbol{\theta})$, which gives the number of companions per star per logarithmic $a$-$M_{c}$ interval. We define $\boldsymbol{\omega}=(\log a,\,\log M_{c})$ for notational compactness and use $\boldsymbol{\theta}$ to denote the vector of model parameters of $\lambda$, which for a histogram model is simply the bin heights.
The hierarchical model likelihood comprises two terms: the probability of detecting $K$ companions with the observed $a$ and $M_{c}$, and the probability of detecting no more companions in the rest of the parameter space. The likelihood for the $j^{\text{th}}$ histogram bin, i.e., two-dimensional interval in $\log a$$\log M_{c}$ space, is
@@
$\mathcal{L}_{j}\propto e^{-\Lambda_{j}}\prod_{k=1}^{K}\Bigg[\frac{1}{N_{k}}\sum_{n=1}^{N_{k}}\frac{\lambda(\boldsymbol{\omega}_{k}^{(n)}|\boldsymbol{\theta})Q(\boldsymbol{\omega}_{k}^{(n)})}{p(\boldsymbol{\omega}_{k}^{(n)}|\boldsymbol{\alpha})}\cdot\mathbbm{1}_{n,k,j}\Bigg]^{w_{k,j}}.$ (1)
@@
In the first term, the Poisson rate parameter $\Lambda_{j}$ gives the expected total number of observed companions among $N_{\star}$ host stars in the $j^{\text{th}}$ bin:
$\Lambda_{j}=N_{*}\int\lambda(\boldsymbol{\omega}|\boldsymbol{\theta})\,Q(\boldsymbol{\omega})\,d\boldsymbol{\omega},$ (2)
@@
where the integral is over the $j^{\text{th}}$ bin and $Q$, the average sensitivity of the CLS to companions as a function of $a$ and $M_{c}$, accounts for missed companions.
@@
The second term gives the probability of detecting $K$ companions with the observed parameters $\{\boldsymbol{\omega}_{k}^{(n)}\}$:
@@
$P(\boldsymbol{\omega}|\boldsymbol{\theta})=\prod_{k=1}^{K}\Bigg[\frac{1}{N_{k}}\sum_{n=1}^{N_{k}}\frac{\lambda(\boldsymbol{\omega}_{k}^{(n)}|\boldsymbol{\theta})Q(\boldsymbol{\omega}_{k}^{(n)})}{p(\boldsymbol{\omega}_{k}^{(n)}|\boldsymbol{\alpha})}\cdot\mathbbm{1}_{n,k,j}\Bigg]^{w_{k,j}}.$ (3)
We use $k$ and $n$ to index the companions and posterior samples, respectively, so that $\boldsymbol{\omega}_{k}^{(n)}$ represents the $n^{\text{th}}$ draw from the posterior of companion $k$. We included $K=191$ companions in our occurrence analysis, and sampled $N_{k}=1000$ draws from each posterior. The indicator function $\mathbbm{1}_{n,k,j}$ equals one when $\boldsymbol{\omega}_{k}^{(n)}$ falls in the $j^{\text{th}}$ bin, and zero otherwise. The interim prior $p$ is the prior we used during orbit fitting: log-uniform in $a$ and $M_{c}$.
Many of the posteriors in our catalog have fractional uncertainties that are comparable to the bin sizes. Following Gilbert et al. (2025), we accounted for these large errors by weighting each companion’s contribution to the likelihood. For each companion/bin pair, we calculated a weight, $w_{k,j}$, equal to the fraction of draws from that companion’s posterior that fell within the bin.
The full likelihood over $J$ bins is
$\mathcal{L}\propto\prod_{j=1}^{J}\mathcal{L}_{j}.$ (4)
We placed uniform priors on the bin heights $\boldsymbol{\theta}$ and sampled our likelihood using emcee.
We repeated our fitting procedure with the extra requirement that the histogram bin heights vary smoothly. We enforced this assumption by drawing bin heights from a Gaussian Process (GP) prior defined by a two-dimensional Matérn–3/2 kernel, with covariance function
$K(\omega_{i},\omega_{j})=\sigma^{2}\left(1+\sqrt{3}d\right)\text{exp}\left(-\sqrt{3}d\right),$ (5)
where $d^{2}=(\frac{\Delta\log a}{\ell_{a}})^{2}+(\frac{\Delta\log M_{c}}{\ell_{M_{c}}})^{2}$ and the covariance scale $\sigma$, the semi-major axis correlation length $\ell_{a}$, and the mass correlation length $\ell_{M_{c}}$ are hyperparameters of the model. We tested both log-normal and log-uniform hyperpriors on these parameters, finding in both cases that the parameters were essentially unconstrained by the data. Given the statistically negligible effect of GP smoothing on our results, we report our non-smoothed results in the remainder of this work.
IV Results
IV.1 Companion occurrence as a function of mass and semi-major axis
@@
We measured occurrence as a function of both mass and separation in 30 cells defined by the following boundaries: $\frac{M_{c}}{M_{\mathrm{Jup}}}=${0.8, 1.6, 3.2, 6.4, 13, 80}, $\frac{a}{AU}=${0.03, 0.1, 0.3, 1, 3, 10, 30}. We highlight our region of interest in the context of the CLS true mass catalog in the right panel of Figure 1. Figure 2 shows the results of these calculations and we provide the occurrence rates in Table 2. The drawback of calculating occurrence in 30 separate bin is that each bin contains few companion detections, yielding large uncertainties on our measurements. We therefore performed a simple model comparison for the semi-major axis distribution of each mass bin: a step function versus a log-linear model, given respectively by
@@
IV.2 The Brown Dwarf Desert at 1-10 AU
@@
Figure 3: Marginalized occurrence rate density of CLS companions between 1–10 AU. Companion occurrence decreases log-linearly with increasing mass up to $\sim$20 $M_{\mathrm{Jup}}$, disfavoring a ‘cut-off’ mass at which companion formation near the ice line abruptly becomes inefficient. We annotate each mass bin with the true occurrence rate. The total occurrence rate between 13–80 $M_{\mathrm{Jup}}$ is $1.1^{+0.5}_{-0.4}\%$.
IV.3 A Rise in Giant Companion Occurrence at the Ice Line
Finally, we examined the occurrence distribution of giant companions as a function of semi-major axis. We used the same separation bins as in Section IV.1, this time marginalizing over companion masses between 0.8–80 $M_{\mathrm{Jup}}$. We found that companion occurrence is approximately constant at close separations, increases abruptly near 1 AU, and remains enhanced to around 10 AU. There is also a marginally significant decrease in occurrence beyond 10 AU, though the uncertainty on our occurrence measurement between 10–30 AU prevents a firm conclusion. Figure 4 shows our marginalized semi-major axis distribution, with the corresponding occurrence rates given in Table 2. Comparing this distribution with those we calculated for separate mass intervals (see Section IV.1), we see that the relatively high occurrence between 1–10 AU exhibited in the low-mass bins contributes to the rise at 1 AU in the marginalized distribution. Meanwhile, the small number of detections between 10–30 AU in all mass intervals results in comparatively low occurrence in this range. We fit this distribution with both a two-level step function and a broken power law. Our broken power law model matches that of Fulton et al. (2021) and is given by
$\lambda_{a}(a|C,\beta,a_{0},\gamma)=C\left(\frac{a}{\text{AU}}\right)^{\beta}\left(1-e^{-(\frac{a}{a_{0}})^{\gamma}}\right),$ (8)
where C is a normalization constant, $\beta$ is the power law index at separations beyond the “break” separation $a_{0}$, and $\beta+\gamma$ is the power law index interior to $a_{0}$. We found that both models performed comparably well, and that the step location reliably fell near 1 AU.
We repeated these calculations using companion minimum mass and modified parameter bounds to match the analysis of Fulton et al. (2021), who calculated the semi-major axis distribution between 0.1–30 AU for companions with $M_{c}\sin i$ =30–6000 $M_{\oplus}$ (0.09–18.9 $M_{\mathrm{Jup}}$). The maximum likelihood parameters of our broken power law fit to this occurrence distribution agreed closely with those derived by Fulton et al. (2021), and we recovered the two key features noted in that study: an abrupt rise in occurrence near 1 AU and a possible fall-off beyond 10 AU. The persistence of the increase in occurrence near 1 AU through changes in binning, mass limits, and $M_{c}$ versus $M_{c}\sin i$ suggests that it is a robust feature of the radial occurrence distribution.
@@ IV.3 A Rise in Giant Companion Occurrence at the Ice Line
V Discussion
V.1 An Astrophysically Motivated Planet-Brown Dwarf Divide
@@
The International Astronomical Union defined brown dwarfs in 2003 as “substellar objects with true masses above the limiting mass for thermonuclear fusion of deuterium $\sim$13 $M_{\mathrm{Jup}}$ ] … no matter how they formed nor where they are located.” [^2] https://web.archive.org/web/20141216075559/http://home.dtm.ciw.edu/users/boss/definition.html This categorization had already been questioned by that time (Burrows et al., 2001), and has faced further scrutiny in the years since (e.g., Spiegel et al. 2011; Schlaufman 2018). Despite the benefit of providing an objective dividing line based on the measurable quantity $M_{c}$, there is no evidence that the processes that form giant planets and brown dwarfs are directly tied to the deuterium burning limit. To better address the nuances of companion formation at the planet-brown dwarf transition, astronomers have searched for definitions motivated by factors such as formation pathway, dynamics, and host star properties. Schlaufman (2018) found evidence for a divide between objects above/below 4–10 $M_{\mathrm{Jup}}$ based on the preference for objects below this range to orbit metal-rich host stars, a trait not exhibited by objects above it. Combining high-contrast imaging and astrometric observations between 5–100 AU, Bowler et al. (2020) found that companions with $M_{c}<15$ $M_{\mathrm{Jup}}$ have preferentially low eccentricities, peaking at $e\lesssim 0.2$, whereas higher-mass companions exhibit a broad range of eccentricities between ee=0.6–0.9. Using a subset of the companion sample of Bowler et al. (2020) but a different framework for calculating parameter priors, Do Ó et al. (2023) found insufficient evidence for a distinction in the eccentricity distributions between low- and high-mass companions. Recent work by our team (Giacalone et al., 2025; Gilbert et al., 2026) examines the role of both metallicity and eccentricity in the planet-BD distinction.
@@
We searched for a similar distinction in the radial occurrence distributions of planets and BDs in Section IV.1. Among the five mass bins we analyzed, we did not find any abrupt change in the radial distributions with increasing mass. The lack of a sharp dividing line in mass among these distributions suggests that “bottom-up” and “top-down” formation mechanisms may produce companions in overlapping mass ranges. Because the majority of our cells contained fewer than five objects in this experiment, we were unable to constrain occurrence at high significance. Thus, some latent separation-dependent trends may exist that will provide a clearer dividing line between planets and brown dwarfs. On the other hand, we also found smooth variation with increasing mass for the marginalized occurrence distribution in Section IV.2, supporting the idea of a “fuzzy” planet-BD boundary.
@@
We found marginal evidence for jumps in occurrence near 1 AU for companions between 0.8–3.2 $M_{\mathrm{Jup}}$. These features likely drive the analogous and more significant jump in the marginalized radial distribution (Section IV.3). Increased core accretion efficiency is expected near the ice line due to the enhancement of icy pebbles and silicate dust that can coagulate via gravitational and/or streaming instabilities in this region (Drażkowska and Alibert, 2017; Hyodo et al., 2019). Thus, the presence or absence of an abrupt increase in occurrence near the ice line may be used to infer whether an object formed by core accretion or gas/disk instability. Increased Doppler baselines, larger stellar samples, and superior orbital fits through the inclusion of Gaia DR4 astrometry (Winn, 2022) will enable occurrence measurements precise enough to confirm or rule out an occurrence jump in narrow mass intervals such as those examined here. Such measurements will place an astrophysically motivated planet-brown dwarf dividing line within reach.
V.2 The Brown Dwarf Desert